NEED THIS IN EXCEL FORM PLEASE

ESPN currently pays the NFL $1.1 billion per year for eight years for the right to exclusively televise Monday Night Football. What is the net present value of this investment if the parent Disnery company has an opportunity interest rate equal to its cost of capital of 9 percent. Fox and CBS agreed to pay $712 million and $622 million respectively for six years to televise Sunday afternoon NFC games. What was that worth?
Answer #1
1- ESPN
Year cash flow present value factor at 9% =1/(1+r)^n r=9% present value of cash flow =cash flow*present value factor
1 1.1 0.917431193 1.009174312
2 1.1 0.841679993 0.925847993
3 1.1 0.77218348 0.849401828
4 1.1 0.708425211 0.779267732
5 1.1 0.649931386 0.714924525
6 1.1 0.596267327 0.65589406
7 1.1 0.547034245 0.601737669
8 1.1 0.50186628 0.552052908
present value of investment sum of present value of cash flow 6.09
2- Fox
Year cash flow present value factor at 9% =1/(1+r)^n r=9% present value of cash flow =cash flow*present value factor
1 712 0.917431193 653.2110092
2 712 0.841679993 599.2761552
3 712 0.77218348 549.7946378
4 712 0.708425211 504.3987503
5 712 0.649931386 462.751147
6 712 0.596267327 424.5423367
present value of investment sum of present value of cash flow 3193.97
3- CBS
Year cash flow present value factor at 9% =1/(1+r)^n r=9% present value of cash flow =cash flow*present value factor
1 622 0.917431193 570.6422018
2 622 0.841679993 523.5249558
3 622 0.77218348 480.2981246
4 622 0.708425211 440.6404813
5 622 0.649931386 404.2573223
6 622 0.596267327 370.8782773
present value of investment sum of present value of cash flow 2790.24